# ChartGalaxy: A Dataset for Infographic Chart Understanding and Generation

### v3

@@ D.3 Example-based Infographic Chart Generation
Appendix E Prompts for Data Processing
E.1 Instruction Dataset for Infographic Chart Understanding
This section details the prompts utilized for constructing our instruction dataset for infographic chart understanding. The prompts generate question-answer pairs for both text-based reasoning and visual-element-based reasoning. These questions cover reasoning categories including Data Identification (DI), Data Comparison (DC), Data Extraction with Condition (DEC), and Fact Checking (FC). Visual-element-based reasoning questions are specifically created by determining if data attributes within a generated question possess corresponding icon representations in the chart image; if such icons exist, the textual attribute is then replaced by its respective icon. In addition to prompt-generated questions, we also incorporate template-based questions from ChartAssistant [39] as a supplement. For Visual Understanding questions, style-related questions (Style Detection, Visual Encoding Analysis, and Chart Classification) are derived from the chart styles used in our generation process and do not require specific prompts.
@@ E.1 Instruction Dataset for Infographic Chart Understanding
E.2 Benchmarking Infographic Chart Code Generation
@@
You are an expert data-visualization engineer and front-end developer. Your task is to take a chart image and generate a HTML file that, when loaded in a browser, reproduces the chart exactly. The chart must be centered in the viewport. # Constraints: No Explanations: Do not include comments, reasoning, or explanatory text. Output only valid HTML with JavaScript code. # Technical Requirements Charting library: Use D3.js to implement the chart. Write the code to be clean, modular, and easy to understand and modify. Single file output: Provide one standalone HTML file that includes everything needed to render the chart. Chart fidelity: Replicate all visual elements—shapes, colors, axes, labels, legends, fonts, line weights, markers—exactly as in the original image. No animations: The chart must render immediately in its final state. Aspect ratio & sizing: The chart’s content area (including margins, paddings, and plot area) must match the original image’s proportions precisely. Image: Recreate any icon/image content using SVG <g> and shapes. Do not use <image>, base64 images, or links. Text: Place all text in <text> leaf nodes. Do not use <tspan> or nested text structures. # Grouping Requirements You should use following class names for SVG groups with specific semantics: title: The title area (may contain title, subtitle, and shapes) image: Each individual icon/image/pictogram (no annotation text, no grouped images) legend: Legend area axis: Axis area # Output Format Return only the following standalone HTML file:
@@
Only output this file. Keep it minimal and strictly within the token limits.
@@
E.3 Example-based Infographic Chart Generation
@@
Appendix F Ethical Considerations and Societal Impacts
Ethical consideration
Our research involves the creation of a large-scale infographic chart dataset, comprising 1,151,087 synthesized charts and 104,519 real charts. We fully respect the intellectual property rights of original infographic designers. The real charts used in our dataset were all sourced from publicly available online platforms, and no proprietary or paywalled content was included. Where applicable, we extracted only the design patterns (chart variations and layout templates), not the original textual or commercial content, in adherence to fair use principles. Our proposed automatic generation method enables scalable generation of high-fidelity infographic-style charts, which can serve as a foundation for future research without infringing upon the rights of original creators.
Our project has been approved by our institution’s internal ethics review. All human subjects involved in our user study have signed the user consent forms and have been provided with fair compensation in accordance with the local minimum wage standards.
Societal impacts
Infographic charts are a powerful medium for communicating complex ideas to broad audiences. Our benchmark and generation pipeline support the development of LVLMs that can better understand and produce such visualizations, enhancing applications in education, journalism, and public communication. By enabling scalable, copyright-respecting data generation, our work lowers barriers for research while promoting ethical development. We believe this contributes meaningfully to improving the visual reasoning capabilities of MLLMs, especially in real-world tasks that require intuitive, data-driven storytelling.

### v5

@@ H.3 Example-based Infographic Chart Generation
Appendix I Prompts for Data Processing
I.1 Instruction Dataset for Infographic Chart Understanding
This section details the construction of our instruction dataset for infographic chart understanding. We categorize the generated question–answer (QA) pairs into two types: prompt-based pairs and template-based pairs.
The prompt-based pairs are produced through LLM prompting. Specifically, Gemini-2.0-Flash is provided with both the tabular data underlying the chart and the corresponding chart image to produce question–answer pairs. These pairs cover two forms of reasoning: text-based reasoning and visual-element-based reasoning. They further span several categories, including Data Identification (DI), Data Comparison (DC), Data Extraction with Condition (DEC), and Fact Checking (FC). For text-based reasoning, the generated questions directly reference the textual data attributes in the chart. For visual-element-based reasoning, questions are created by first determining whether the data attributes referenced in a generated question correspond to icon representations in the chart image; when such icons are present, the textual attributes are replaced by their respective icons.
The template-based pairs are mainly derived by adapting templates from ChartAssistant (Meng et al., 2024). In addition, style-related questions for Visual Understanding (e.g., Style Detection, Visual Encoding Analysis, and Chart Classification) are also template-based, since they are derived directly from the chart styles used in our generation process and do not rely on LLM prompting.
@@ I.1 Instruction Dataset for Infographic Chart Understanding
I.2 Benchmarking Infographic Chart Code Generation
@@
You are an expert data‐visualization engineer and front‐end developer. Your task is to take a chart image and generate a HTML file that, when loaded in a browser, reproduces the chart exactly. The chart must be centered in the viewport. # Constraints: No Explanations: Do not include comments, reasoning, or explanatory text. Output only valid HTML with JavaScript code. # Technical Requirements Charting library: Use D3.js to implement the chart. Write the code to be clean, modular, and easy to understand and modify. Single file output: Provide one standalone HTML file that includes everything needed to render the chart. Chart fidelity: Replicate all visual elements—shapes, colors, axes, labels, legends, fonts, line weights, markers—exactly as in the original image. No animations: The chart must render immediately in its final state. Aspect ratio & sizing: The chart’s content area (including margins, paddings, and plot area) must match the original image’s proportions precisely. Image: Recreate any icon/image content using SVG <g> and shapes. Do not use <image>, base64 images, or links. Text: Place all text in <text> nodes. Do not use <tspan> or nested text structures. # Grouping Requirements You should use following class names for SVG groups with specific semantics: title: The title area (may contain title, subtitle, and shapes) image: Each individual icon/image/pictogram (no annotation text, no grouped images) legend: Legend area axis: Axis area # Output Format Return only the following standalone HTML file:
@@
Only output this file. No comments or explanation. Keep it minimal and strictly within the token limits.
@@
I.3 Example-based Infographic Chart Generation
@@
Appendix J The use of LLMs
LLMs were used solely to aid or polish writing; they were not used for research ideation, analysis, or any other part of the manuscript.